### 15 Apr 2026 v1

# Enhancing LLM-based Search Agents via Contribution Weighted Group Relative Policy Optimization

Abstract
Search agents extend Large Language Models (LLMs) beyond static parametric knowledge by enabling access to up-to-date and long-tail information unavailable during pretraining. While reinforcement learning has been widely adopted for training such agents, existing approaches face key limitations: process supervision often suffers from unstable value estimation, whereas outcome supervision struggles with credit assignment due to sparse, trajectory-level rewards. To bridge this gap, we propose Contribution-Weighted GRPO (CW-GRPO), a framework that integrates process supervision into group relative policy optimization. Instead of directly optimizing process rewards, CW-GRPO employs an LLM judge to assess the retrieval utility and reasoning correctness at each search round, producing per-round contribution scores. These scores are used to rescale outcome-based advantages along the trajectory, enabling fine-grained credit assignment without sacrificing optimization stability. Experiments on multiple knowledge-intensive benchmarks show that CW-GRPO outperforms standard GRPO by 5.0% on Qwen3-8B and 6.3% on Qwen3-1.7B, leading to more effective search behaviors. Additional analysis reveals that successful trajectories exhibit concentrated contributions across rounds, providing empirical insight into search agent tasks.
*Equal contribution. Corresponding authors.1Our code is available at https://github.com/zsxmwjz/CW-GRPO.
Figure 1: CW-GRPO assesses the quality of each search round within successful trajectories and to concentrate the learning advantage on high-quality search rounds, thereby enabling round-level credit assignment.
1 Introduction
Search agents empower large language models (LLMs) to move beyond fixed parametric knowledge by providing access to real-time information and specialized facts absent from pre-training [li2025searcho1agenticsearchenhancedlarge, zhang2025deep]. Through iterative retrieval of external evidence and integration into the reasoning process, search agents ground model predictions in verifiable sources and substantially enhance the factual reliability of LLM-based systems [zhao2026retrieval, li2025matching, zhou2024trustworthiness, wang2025actingreasoningmoreteaching]. Consequently, they have become a key role in enabling LLMs to perform knowledge-intensive and fact-sensitive tasks [glass2022re2g, sawarkar2024blended].
While reinforcement learning (RL) offers a significant paradigm for training such agents, current RL approaches suffer from fundamental limitations [lee2024multi]. Depending on the level of supervision, they can be broadly categorized into process supervision and outcome supervision [uesato2022solving]. Process supervision [lightman2023letsverifystepstep, zhu2025chain, wang2025stepsearchignitingllmssearch] assigns round-level rewards and is commonly optimized with actor–critic methods such as Proximal Policy Optimization [schulman2017proximalpolicyoptimizationalgorithms], but learning reliable critics over diverse intermediate states is unstable and frequently leads to biased advantages and brittle training [gurumurthy2023practical, hu2025reinforcestabilizingcriticfreepolicy]. Outcome supervision, by contrast, relies solely on final answer correctness, which makes it difficult to attribute task success to individual search rounds, leading to a credit assignment problem and obscuring the uneven contributions of intermediate decisions. Recent advances, such as Group Relative Policy Optimization [shao2024deepseekmathpushinglimitsmathematical], improve training stability and memory efficiency, making outcome-supervised optimization more practical; however, they[jin2025searchr1trainingllmsreason, li2025searcho1agenticsearchenhancedlarge, wang2025actingreasoningmoreteaching, li2025r3] do not alter the sparsity of the reward signal, and the credit assignment issue remains unresolved.
To address the aforementioned challenges, we propose Contribution-Weighted GRPO (CW-GRPO), which reformulates process supervision as a problem of modulating outcome-derived advantages, rather than directly optimizing process rewards. Specifically, CW-GRPO retains the stability of the standard GRPO by computing outcome advantages through group-relative comparisons. We use an LLM judge to assess each search round based on retrieval utility and reasoning correctness. These metrics are synthesized into a contribution score, based on which serves as a scaling factor to redistribute the outcome advantage across the trajectory. By avoiding direct comparison of process rewards, CW-GRPO incorporates process supervision into the GRPO framework, amplifying learning signals for informative rounds while suppressing those for redundant or unproductive ones, as illustrated in Figure 1.
Empirical results across multiple knowledge-intensive benchmarks demonstrate that CW-GRPO consistently outperforms both outcome-supervised and process-supervised baselines, enabling more effective search behaviors. Specifically, CW-GRPO achieves significant performance gains over standard GRPO, with relative improvements of $5.0\%$ on Qwen3-8B and $6.3\%$ on Qwen3-1.7B. The contributions of this work are summarized as follows:
  • A Reframing of Process Supervision: We introduce a new perspective that treats process supervision for search agents as advantage reallocation guided by round contribution, rather than explicit process reward estimation.
  • Contribution-Weighted GRPO: We propose a new optimization framework that brings process-level credit assignment into GRPO without requiring unstable value functions.
  • Empirical Characterization of Search Contribution: We provide empirical evidence showing a structural characteristic in search agent tasks: the contribution to task success is highly concentrated in informative rounds rather than being uniformly distributed.
2 Related Work
2.0.0.1 Agentic Search.
Large language models encode knowledge in static parameters, limiting access to up-to-date and long-tail knowledge. Retrieval-augmented generation (RAG) addresses this by enabling models to access external knowledge sources [zhang2026opennoveltyllmpoweredagenticverifiable, zhao2026retrieval, li2025matching, zhou2024trustworthiness]. An early paradigm is search-as-a-tool, where LLMs are prompted or fine-tuned to interact with static corpora. IRCoT [trivedi2023interleavingretrievalchainofthoughtreasoning] and FLARE [zhang2025flarefeedforwardgeometryappearance] use prompts to guide iterative reasoning and relieve model uncertainty, and Self-RAG [asai2023selfraglearningretrievegenerate] adopts supervised fine-tuning with reflection tokens for adaptive retrieval. More recently, agentic search enables LLMs to interact with real-world search engines for dynamic information access. WebGPT [nakano2022webgptbrowserassistedquestionansweringhuman] pioneers this direction using RLHF. Search-o1 [li2025searcho1agenticsearchenhancedlarge] integrates agentic search into large reasoning models via prompting, while Search-R1 [jin2025searchr1trainingllmsreason] and R1-Searcher [song2025r1searcherincentivizingsearchcapability] adopt RL for end-to-end training with outcome-based rewards. However, outcome supervision assigns uniform credit across all search rounds, failing to distinguish pivotal searches from redundant ones.
2.0.0.2 Outcome Supervision and Process Supervision.
A key challenge in training search agents lies in credit assignment across multi-round search trajectories. Outcome supervision, widely adopted in RL methods like PPO [schulman2017proximalpolicyoptimizationalgorithms] and GRPO [shao2024deepseekmathpushinglimitsmathematical], assigns rewards only based on the final answer, treating all immediate rounds equally and ignoring their unequal contributions. Process supervision addresses this via round-level feedback for finer-grained credit assignment. Early approaches to process supervision rely on learned Process Reward Models (PRMs), but often suffer from costly round-level annotation requirements and limited out-of-domain generalizability [lightman2023letsverifystepstep, huang-etal-2025-empirical]. To avoid training explicit reward models, recent works increasingly leverage frozen LLMs as external evaluators to assess intermediate rounds. These methods use LLM-based judges to assess reasoning quality and then assign round-level rewards, enabling process supervision without training dedicated PRMs [zhang2025criticsearchfinegrainedcreditassignment]. These approaches typically optimize judge-produced signals directly as rewards, making them sensitive to evaluation noise and calibration errors. In contrast, CW-GRPO uses LLM-derived signals only to redistribute outcome-based advantages, avoiding direct optimization over noisy intermediate rewards and thereby enabling more stable training.
3 Contribution-Weighted GRPO
As motivated in Section 1, current search agents suffer from the credit assignment problem where sparse outcome rewards cannot distinguish between pivotal and redundant search rounds. To address this, we propose Contribution-Weighted Group Relative Policy Optimization (CW-GRPO). Unlike traditional process supervision that requires learning a fallible critic, CW-GRPO reframes process-level signals as dynamic scaling factors that redistribute trajectory-level advantages. The overall framework and its core mechanism of round-level credit assignment are presented in Figure 2.
3.1 Task Formulation
We consider a search agent policy $\pi_{\theta}$ tasked with answering a question q. The agent interacts with an environment (search engine) through a trajectory $\tau$ of $T$ rounds:
$\small\tau=\big((s^{1},a^{1}),(s^{2},a^{2}),\dots,(s^{T},a^{T})\big),$ (1)
where for each round $t<T$, the action $a^{t}$ consists of a reasoning chain followed by an invocation of the search tool. The environment then returns retrieved documents to form the state $s^{t+1}$. The last action $a^{T}$ is the final answer.
Training proceeds in groups. For each question, we sample a group of $G$ trajectories $\{\tau_{i}\}_{i=1}^{G}$ under the current policy. Each trajectory $\tau_{i}$ receives a scalar outcome reward $R_{i}$ based on the exact match (EM) of its final answer. As in GRPO, these outcome rewards are only used for relative comparison within the group.
Figure 2: CW-GRPO employs an LLM judge to assess both the retrieval utility and the reasoning correctness at each individual search round. Based on these two signals, it computes a contribution factor for every search round, which is subsequently applied to the original advantage obtained from group-relative comparisons, thereby producing the final round advantage. As a result, training signals from highly contributive rounds are amplified, while those from low-contribution rounds are attenuated.
3.2 Outcome-Level Advantage
Following the GRPO framework, we first establish a baseline for performance by comparing trajectories within the same group. This avoids the need for a value-function critic. For each trajectory $\tau_{i}$, we compute a normalized outcome advantage:
$\small A_{i}^{O}=\frac{R_{i}-\text{mean}\{R_{i}\}|_{i=1}^{G}}{\text{std}\{R_{i}\}|_{i=1}^{G}}.$ (2)
While $A_{i}^{O}$ effectively identifies which trajectories are better than others, it remains temporally coarse, assigning the same credit to every search round within a successful trajectory, regardless of whether a specific round was informative or distracting.
3.3 Round-Level Contribution Estimation
The effectiveness of CW-GRPO fundamentally depends on accurately identifying pivotal rounds within a search trajectory. Accordingly, we decompose the quality of each round into two orthogonal binary signals: retrieval utility and reasoning correctness. Crucially, we adopt a conjunctive formulation, under which a round is considered contributive only if it satisfies both criteria, namely retrieving novel, task-relevant evidence and maintaining logically consistent reasoning. This design reflects an inherent structural property of multi-round search: effective task solving requires not only the acquisition of useful information but also its correct interpretation in context. Retrieval without sound reasoning may lead to incorrect interpretation or misuse of useful evidence; conversely, correct reasoning without informative retrieval fails to make meaningful progress toward solving the task. The conjunctive gating mechanism thus serves as a conservative filter, isolating rounds whose contributions are causally aligned with task success, thereby providing a low-noise and more reliable signal for advantage reallocation.
3.3.1 Signal Definition and Rubrics
To ensure the robustness of the LLM-derived signals, we employ a rubric-based assessment where an LLM judge is guided by a set of predefined criteria. For each search round $t<T_{i}$ in trajectory $\tau_{i}$, the values of the two signals are determined as follows:
  • Retrieval utility $u_{i}^{t}\in\{0,1\}$: This signal assesses the information gained from external sources. A round is assigned $u_{i}^{t}=1$ only if the retrieved documents contain novel, task-relevant evidence that was not present in the context of previous rounds. This prevents the agent from being rewarded for redundant or circular retrieval queries.
  • Reasoning correctness $v_{i}^{t}\in\{0,1\}$: This signal estimates the internal consistency of the agent. A round is assigned $v_{i}^{t}=1$ if the reasoning chain correctly interprets the current context and maintains a logical path toward the final answer. This ensures that the agent is not getting the right answer for the wrong reasons.
The conjunctive contribution signal is defined as the logical product $p_{i}^{t}=u_{i}^{t}\cdot v_{i}^{t}$. By enforcing a binary bottleneck, we treat each search round as a discrete “gate”—either it contributes a necessary building block to the solution, or it does not.
3.3.2 Judge Calibration and Reliability
To ensure that the LLM judge aligns with expert intuition, we conducted a rigorous calibration process. We manually annotated a subset of trajectories (comprising 97 distinct search rounds) to serve as a gold standard.
The rubrics were iteratively refined to minimize ambiguity. In our final setup, the LLM judge achieved a 95% consensus rate with human expert annotations on both retrieval utility and reasoning correctness. This high level of agreement provides a reliable foundation for the subsequent advantage redistribution, ensuring that policy gradients are directed by high-fidelity process signals.
3.4 Adaptive Weighting Mechanism
We then convert these raw signals into a normalized contribution weight $c_{i}^{t}$. Importantly, we treat successful and failed trajectories differently, as a robustness-oriented design.
For successful trajectories ($R_{i}=1$), rounds that introduce new task-relevant evidence and maintain logically consistent reasoning are typically aligned with the final successful outcome. This enables reliable credit redistribution across rounds. A temperature-controlled softmax is therefore adopted to emphasize high-contribution rounds:
$\small c_{i}^{t}=\frac{\exp\big(\alpha p_{i}^{t}\big)}{\sum_{t^{\prime}=1}^{T_{i}-1}\exp\big(\alpha p_{i}^{t^{\prime}}\big)},\quad t<T_{i},\text{ if }R_{i}=1,$ (3)
where $\alpha$ is a hyperparameter controlling the "sharpness" of the redistribution. A higher $\alpha$ forces the model to learn primarily from the most pivotal rounds.
For failed trajectories ($R_{i}=0$), round-level attribution is substantially more ambiguous. Many intermediate rounds exhibit reasonable behavior, such as attempting to retrieve missing evidence or issuing follow-up queries for verification. However, these rounds may still fail to obtain useful information because the required knowledge is not covered by the corpus or cannot be retrieved due to limitations of the retriever. In such cases, the failure does not reflect an identifiable error in the agent’s decision at a specific round. As a result, assigning differentiated weights based on round-level signals is unreliable and can introduce spurious supervision. A more detailed analysis of such failure patterns is provided in Appendix A. To mitigate this issue, we assign a uniform contribution across all rounds:
$\small c_{i}^{t}=\frac{1}{T_{i}-1},\quad t<T_{i},\text{ if }R_{i}=0.$ (4)
This design preserves the stability of outcome-based learning while avoiding noisy round-level credit assignment under ambiguous attribution. Failed trajectories are still utilized through outcome-level comparison, and credit redistribution is applied only when attribution is reliable.
By construction, $\sum_{t=1}^{T_{i}-1}c_{i}^{t}=1$.
3.5 Advantage Reallocation and Optimization
The core of CW-GRPO is the reallocated advantage $A_{i}^{t}$. For all search rounds $t<T_{i}$ of trajectory $\tau_{i}$, we scale the outcome advantage by the normalized contribution. And for the final answer round $t=T_{i}$, we directly use the outcome advantage:
$\small A_{i}^{t}=A_{i}^{O}\cdot c_{i}^{t}\cdot(T_{i}-1),\text{ if }t<T_{i};\quad A_{i}^{T_{i}}=A_{i}^{O}.$ (5)
The scaling factor $(T_{i}-1)$ ensures the magnitude of the total learning signal for the trajectory remains conserved, i.e.,
$\small\frac{1}{(T_{i}-1)}\sum_{t=1}^{T_{i}-1}A_{i}^{t}=A_{i}^{O}.$ (6)
Finally, we optimize the policy using the clipped surrogate objective:
$\small\mathcal{L}(\theta)=-\mathbb{E}\Big[\min\big(rA,\text{clip}(r,1-\epsilon,1+\epsilon)A\big)\Big],$ (7)
where $\epsilon$ is a hyper-parameter for clipping, $r=\frac{\pi_{\theta}(a\mid s)}{\pi_{\theta_{old}}(a\mid s)}=\exp(\log\pi_{\theta}-\log\pi_{\theta_{old}})$ is the importance sampling ratio.
In summary, CW-GRPO achieves a seamless integration of process supervision into the GRPO framework. Our method achieves effective credit assignment across iterative search rounds by reallocating trajectory-level advantages based on each round’s contribution, rather than estimating absolute intermediate rewards. This design preserves the training stability of group-relative optimization while ensuring the policy learns more effectively from the specific decisions that drive task success.
4 Setup
4.1 Datasets and Evaluation Metrics
We evaluate our method on two categories of knowledge-intensive tasks: (1) General Question Answering: Natural Questions (NQ) [kwiatkowski-etal-2019-natural], TriviaQA [joshi2017triviaqalargescaledistantly], and PopQA [mallen2023trustlanguagemodelsinvestigating]. (2) Multi-Hop Question Answering: HotpotQA [yang2018hotpotqadatasetdiverseexplainable], 2WikiMultiHopQA [ho2020constructingmultihopqadataset], Musique [trivedi2022musiquemultihopquestionssinglehop], and Bamboogle [press2023measuring]. We merge the training sets of NQ and HotpotQA as the training data. To mitigate the impact of decoding variance, we use Avg@4 Exact Match (EM) as the evaluation metric, averaging the EM scores over four sampled responses per task.
4.1.0.1 Hard-Case Evaluation Set.
In standard settings, models often bypass external search by leveraging internal parametric knowledge, which obscures their true agentic search proficiency. In light of this, we adopt a more challenging protocol using AgentGym-SearchQA-test, a 400-sample test set from AgentGym-RL [xi2025agentgymrltrainingllmagents]. These samples are specifically filtered from cases where the Qwen2.5-72B-Instruct [yang2024qwen25mathtechnicalreportmathematical] model failed, representing the "hardest" distribution for current LLMs. This setup shifts the focus from simple parametric recall to complex information-seeking and reasoning capabilities, explicitly emphasizing whether agents retrieve and utilize external evidence rather than relying on parametric knowledge.
4.1.0.2 Strict Retrieval Constraint.
To further emphasize the model’s search agentic capabilities, we include a specific instruction in the system prompt: the model is prohibited from using any parametric knowledge beyond basic common sense.
4.2 Baselines
We compare our approach against two categories of search agents: (1) Outcome-Supervised Methods: Search-R1-PPO and Search-R1-GRPO [jin2025searchr1trainingllmsreason], which optimize the final answer via group relative policy optimization. (2) Process-Supervised Methods: R3-RAG [li2025r3] and MT-PPO [wei2025reinforcingmultiturnreasoningllm], which utilize explicit process rewards and PPO frameworks.
We also include results from state-of-the-art closed-source models (GPT-5.1 [singh2025openaigpt5card], Gemini-3-Pro-Preview [google2025gemini3pro], Qwen3-Max [qwen3max]) as strong upper-bound references, alongside powerful open-source models (DeepSeek-V3.2 [deepseekai2025deepseekv32pushingfrontieropen], Qwen3-32B [qwen3technicalreport]) for comparison.
4.2.0.1 Note on Comparison:
Because of our uniquely challenging test set and the strict use of a no-parametric-knowledge prompt, the baseline performance reported in this paper differs from that reported in the original publications.
4.3 Implementation Details
We use Qwen3-8B and Qwen3-1.7B [qwen3technicalreport] as base models due to their tool-calling capabilities. Training is conducted using the veRL [sheng2024hybridflow] framework. The maximum context length is set to 9192 tokens, with up to 10 interaction rounds. We employ SGLang [zheng2024sglang] as the inference engine for rollout, setting both temperature and top_p both to 1.0.
Parameter GRPO / CW-GRPO PPO-based (Actor) PPO-based (Critic)
Learning Rate $\times 10^{-6} 1×1$$\times 10^{-6}$ $1\times 10^{-5}$
Warmup Ratio 0.285 0.285 0.015
Training Steps 200 200 200
Batch Size 32 128 128
Group Size 4 1 1
GAE $\gamma/\lambda$ - 1.0/1.0 1.0/1.0
Table 1: Detailed training hyperparameters of CW-GRPO and baselines.
4.3.0.1 Retrieval Infrastructure.
Following Search-R1, we use the 2018 Wikipedia dump [wikidump] as the knowledge source. We employ E5 [wang2022text] as the retriever, with the number of retrieved documents $k$ set to 3.
4.3.0.2 Training Hyperparameters.
For all RL-based experiments, we use a KL-divergence regularization coefficient $\beta=0.001$ and a clipping ratio $\epsilon=0.2$. Specifically, for our proposed CW-GRPO, we apply GPT-oss-120B [openai2025gptoss120bgptoss20bmodel] as the judge model and set $\alpha=\infty$. This configuration ensures that for trajectories resulting in task success, the learning signals during search iterations are exclusively concentrated on those rounds where the conjunctive contribution signal is $1$. Other specific parameters are detailed in Table 1. The differences in hyperparameter configurations between GRPO-based and PPO-based methods stem from their distinct algorithmic characteristics (e.g., GRPO relies on intra-group relative comparisons while PPO optimizes independent trajectories), and we ensure fairness by normalizing the total number of sampled trajectories per training step.
5 Experimental Results
Method General QA Multi-Hop QA Overall
NQ TriviaQA* PopQA* HotpotQA 2wiki* Musique* Bamboogle*
Proprietary Models
GPT-5.1 31.0 60.0 38.5 36.0 43.5 13.8 39.0 34.44
Gemini-3-Pro-Preview 31.0 66.0 32.5 42.0 38.5 14.0 42.0 35.00
Qwen3-Max 31.5 62.5 30.5 39.5 49.0 18.5 42.5 36.56
Open-sourced Models
DeepSeek-V3.2-Thinking 24.5 47.0 22.0 30.0 24.0 12.25 29.5 25.19
Qwen3-32B 20.5 53.0 22.0 28.0 33.5 8.5 30.0 25.50
Qwen3-1.7B 13.5 34.5 18.0 11.0 6.0 2.5 4.5 11.56
Search-R1-PPO 19.5 50.5 25.0 25.5 24.5 5.0 14.0 21.06
Search-R1-GRPO 18.0 45.5 25.0 27.0 22.5 6.75 16.5 21.00
R3-RAG 20.0 47.5 25.5 22.0 18.5 5.25 13.5 19.69
MT-PPO 25.5 49.0 25.0 24.0 31.5 6.0 13.0 22.19
CW-GRPO 22.0 47.5 25.0 24.0 28.5 7.0 17.5 22.31
Qwen3-8B 23.5 53.5 22.5 27.5 27.0 10.0 30.0 25.50
Search-R1-PPO 29.5 57.0 27.5 33.5 36.0 12.3 33.5 30.19
Search-R1-GRPO 33.5 57.0 26.0 33.5 32.5 12.5 31.5 29.88
R3-RAG 26.5 57.5 26.0 32.0 32.0 11.3 33.5 28.75
MT-PPO 25.5 56.5 26.0 32.0 37.0 12.5 31.5 29.19
CW-GRPO 35.5 57.5 28.5 33.5 33.0 13.0 37.0 31.38
Table 2: Avg@4 Exact Match on seven subsets of AgentGym-SearchQA-test. † / * indicate in-domain and out-of-domain datasets. CW-GRPO achieves the best performance among methods built on the same backbone.
5.1 Main Results
Table 2 presents the main experimental results on hard knowledge-intensive QA benchmarks under no-parametric-knowledge constraints.
5.1.0.1 Overall Performance.
Our experimental results demonstrate the robust superiority of CW-GRPO, which achieves state-of-the-art performance across both model scales, outperforming all outcome- and process-supervised baselines on the identical backbone. On Qwen3-8B, CW-GRPO achieves an overall score of 31.38, representing a $5.0\%$ relative improvement over Search-R1-GRPO (29.88). On Qwen3-1.7B, CW-GRPO improves the overall performance from 21.00 to 22.31, corresponding to a $6.3\%$ relative gain. These results indicate that CW-GRPO scales favorably across model sizes and is particularly effective in low-capacity regimes, where efficient search credit assignment is critical.
5.1.0.2 Comparison with Open- and Closed-Source Models.
With Qwen3-8B as the backbone, CW-GRPO surpasses all evaluated open-source models, including DeepSeek-V3.2-Thinking and Qwen3-32B, despite operating under a stricter evaluation setting. However, a clear performance gap remains between CW-GRPO and large closed-source models such as GPT-5.1, Gemini-3-Pro-Preview, and Qwen3-Max, suggesting that stronger base models and proprietary training data still provide advantages under this challenging hard-case distribution.
5.1.0.3 General QA vs. Multi-Hop QA.
CW-GRPO demonstrates consistent improvements on both General QA and Multi-Hop QA tasks compared to outcome-supervised and process-supervised baselines. On Multi-Hop QA, CW-GRPO yields the most pronounced gains. On both Qwen3-1.7B and Qwen3-8B, it improves performance on HotpotQA, Musique, and Bamboogle simultaneously, indicating stronger long-horizon reasoning and evidence aggregation. On General QA, CW-GRPO remains competitive across datasets. While it achieves the best overall General QA performance on Qwen3-8B, its improvement on Qwen3-1.7B is smaller than that of MT-PPO. This suggests that for simpler single-hop questions, dense process supervision can still be effective for very small models, whereas CW-GRPO shows clearer advantages as reasoning depth increases.
5.1.0.4 Summary.
Overall, these results demonstrate that CW-GRPO provides a robust and scalable improvement over existing outcome- and process-supervised search agents, particularly for multi-hop reasoning under strict retrieval constraints. The consistent gains across model sizes highlight the effectiveness of contribution-weighted optimization in guiding search behavior, even in extremely challenging evaluation settings.
5.2 Sharpness of Advantage Reallocation
Based on Qwen3-8B, we analyze the effect of the sharpness parameter $\alpha$, which controls how outcome advantages are redistributed across search rounds based on their estimated contributions. Setting $\alpha=0$ yields uniform weighting over the search rounds and reduces the method to standard GRPO, whereas $\alpha=\infty$ corresponds to CW-GRPO, where the advantage is fully concentrated on rounds with both retrieval utility and correct reasoning.
As shown in Table 3, increasing $\alpha$ generally leads to improved overall performance, with the best overall result achieved at $\alpha=\infty$. This trend suggests that, in successful search trajectories, the contribution to task success is highly uneven across rounds. Rather than accumulating gradually, effective progress is driven by high-quality search rounds that exhibit retrieval utility and correct reasoning. For search agent tasks, effective credit assignment requires acknowledging this highly concentrated contribution structure.
5.3 Supervision Decomposition
We ablate the two supervision signals in CW-GRPO by separately removing retrieval utility or reasoning correctness supervision while keeping all other components unchanged.
According to Table 4, removing either signal consistently degrades performance across both single-hop and multi-hop QA tasks, suggesting that neither signal alone is sufficient for reliable credit assignment during search.
Without retrieval utility supervision, rounds with coherent but uninformative reasoning can receive positive credit. Without reasoning correctness supervision, rounds can be assigned positive contribution weights for retrieving relevant documents even when they are incorrectly interpreted, weakening the link between evidence acquisition and correct reasoning. In either setting, contribution weights may be assigned to rounds that do not make a substantive contribution to task completion, introducing noise into advantage reallocation and harming overall performance.
   $\alpha$ General QA Multi-Hop QA Overall
NQ TriviaQA PopQA HotpotQA 2wiki Musique Bamboogle
0 33.5 57.0 26.0 33.5 32.5 12.5 31.5 29.88
1 27.0 58.0 28.0 32.5 35.0 12.8 31.5 29.69
3 33.0 58.5 31.0 29.5 36.0 13.0 31.0 30.63
5 27.0 58.5 27.0 35.0 39.0 12.0 33.5 30.50
$\infty$ 35.5 57.5 28.5 33.5 33.0 13.0 37.0 31.38
Table 3: Effect of the sharpness parameter $\alpha$ on search agent performance. Larger $\alpha$ value generally yields better overall accuracy, indicating that concentrating learning signals on high-contribution search rounds is beneficial.
Method General QA Multi-Hop QA Overall
NQ TriviaQA PopQA HotpotQA 2wiki Musique Bamboogle
CW-GRPO 35.5 57.5 28.5 33.5 33.0 13.0 37.0 31.38
wo-retrieval 27.0 53.5 26.0 31.0 37.5 12.0 34.5 29.19
wo-reasoning 23.5 55.5 31.0 34.0 28.5 12.25 31.5 28.56
Table 4: Ablation on contribution supervision in CW-GRPO. We remove supervision on retrieval utility (wo-retrieval) or reasoning correctness (wo-reasoning), allowing a search round to contribute if only one condition is satisfied. Dropping either signal leads to inferior performance compared to full CW-GRPO, demonstrating that both are necessary for effective credit assignment in search agents.
5.4 Training Dynamics
As shown in Figure 3, the optimization remains stable across training, with no signs of divergence or collapse. The Exact Match score steadily improves, while the number of task rounds stays well-controlled and converges to a stable range. In addition, both retrieval utility and reasoning correctness exhibit consistent upward trends, indicating progressive improvement in the agent’s search behavior. These results show that CW-GRPO achieves stable optimization while improving both final accuracy and the quality of the search process.
We further analyze the effect of the conjunctive gate mechanism during training. Among all search rounds in successful trajectories, approximately 13.8% are filtered out (i.e., assigned zero contribution), indicating that the gate meaningfully reshapes the distribution of learning signals rather than acting as a near-identity transformation.
5.5 Case Study
Figure 4 compares the search trajectories of Qwen3-8B before and after CW-GRPO training on a factual question regarding John I. Jenkins’s tenure start. The untrained model issues only a single query fails to distinguish between the election year and the actual commencement of his term, leading to an incorrect answer based on this temporal confusion.
In contrast, the CW-GRPO trained model adopts a qualitatively different strategy by issuing multiple parallel queries in the first round. This expanded search coverage allows the agent to recall a broader set of documents, including explicit evidence that Jenkins’s term began in 2005. This case illustrates that CW-GRPO not only improves retrieval utility through parallelization but also strengthens the agent’s resistance to distractor evidence. By learning to prioritize direct answer-bearing evidence over superficially relevant but misleading signals, the model achieves more robust and grounded search-based decision making.
Figure 3: Training curves of CW-GRPO on Qwen3-8B. We report the exponential moving averages of Exact Match, retrieval utility, and reasoning correctness over training steps.
Figure 4: Case study of CW-GRPO on Qwen3-8B.
6 Discussion: Generalization Across Tasks and Models
CW-GRPO is motivated by the general principle of redistributing outcome-derived advantages according to per-round contribution estimates. This principle is broadly applicable beyond the specific experimental settings considered in this work.
6.0.0.1 Task generalization.
CW-GRPO can be applied to multi-round agentic tasks where the contributions of intermediate rounds can be reasonably estimated. In long-horizon decision-making environments(e.g., Webshop [yao2022webshop]), individual actions should correspond to incremental progress toward latent subgoals. This enables a judge to assess state improvements at each round and facilitates per-round credit assignment.
The framework also extends to single-round but structurally decomposable reasoning tasks. In settings such as mathematical reasoning(e.g., MATH [hendrycks2021measuringmathematicalproblemsolving]), model outputs typically consist of multi-step derivations. When a trajectory fails, the error can often be localized to specific steps; these steps can then be assigned higher responsibility for failure, enabling a CW-GRPO variant based on step-level attribution.
However, CW-GRPO is less suitable for tasks where per-step contribution is ill-defined or inherently unstable. This includes highly subjective generation tasks (e.g., open-ended creative writing) and perceptual generation tasks (e.g., image synthesis), where holistic judgments dominate and assigning reliable credit to intermediate steps via low-variance signals is challenging.
6.0.0.2 Model generalization.
CW-GRPO is model-agnostic and does not rely on a specific backbone architecture. Nevertheless, stronger base models tend to be more sensitive to the accuracy of per-round supervision, placing higher demands on the quality of the judge or rubric design, which may limit the practical gains of the method.
7 Conclusion
In this work, we propose CW-GRPO, which reallocates outcome advantages according to round-level contribution rather than explicit process rewards to tackle the credit assignment challenge in training LLM-based search agents. CW-GRPO retains GRPO’s stability while enabling fine-grained credit assignment along search rounds. Our evaluations show that CW-GRPO consistently outperforms baselines, and task success is closely associated with the high-quality rounds, highlighting the importance of modeling contribution concentration in search trajectories. Overall, CW-GRPO offers a simple and effective way to incorporate process-level insights into outcome-supervised training for scalable agentic systems.
Limitations
Despite the advances demonstrated by CW-GRPO in alleviating credit assignment under outcome supervision, several limitations remain.
First, CW-GRPO reallocates advantages only for successful trajectories. For failed trajectories, it does not perform fine-grained credit assignment, as attributing task failure to specific search rounds remains challenging. Consequently, CW-GRPO lacks explicit modeling of which intermediate decisions contribute to failure, leaving failure cases underutilized for process-aware learning.
Second, CW-GRPO adopts a conjunctive gating mechanism for modeling round-level contributions, which improves robustness to judging noise but limits the expressiveness of contribution representation. In addition, the framework operates at the round level and does not naturally extend to token-level supervision, making finer-grained credit assignment beyond rounds hard to incorporate.
Finally, CW-GRPO relies on an external LLM judge for contribution estimation. While we find the associated computational cost to be manageable in our setting (see Appendix C), this design still introduces additional inference cost and dependency on judge quality. Exploring lighter-weight supervision signals (e.g., distilled or heuristic judges) may further improve scalability, especially in more resource-constrained or latency-sensitive scenarios.
Appendix
Appendix A Analysis of Failure Trajectories and Credit Assignment Ambiguity
In this section, we discuss why round-level credit assignment is unreliable in failed trajectories, complementing the design choice discussed in Section 3.4. The key challenge is that failed trajectories typically lack clearly identifiable low-quality rounds that can be reliably attributed to the final failure. Such ambiguity leads to a high risk of introducing noise when selectively reweighting search rounds. Concretely:
  • Redundant/verification queries: These rounds look redundant but may serve an implicit verification role (re-checking ambiguous evidence). Heuristically marking such rounds as low-quality would risk removing legitimate verification behavior.
  • Cascading Errors from Misinterpreted Evidence: Later retrieval rounds may be misled by earlier misinterpretations of superficially relevant yet non-supportive documents, and therefore should not be regarded as the primary source of failure.
  • Attempts that failed to obtain key information: When a search round recalls no useful documents, it usually shows no obvious error — the failure is caused by insufficient corpus coverage or inadequate retriever recall, not a clearly “bad” decision the agent made.
  • Premature overconfidence: This usually occurs at the answer round, not during intermediate search rounds, and thus falls outside the scope of round-level search credit assignment.
A failure pattern that yields a reliably detectable low-quality signal is similar-entity confusion. Because most failure patterns are inherently ambiguous at round granularity, treating failed trajectories differently (uniform contribution) avoids reinforcing spurious signals and preserves GRPO’s stable outcome-based baseline.
Appendix B LLM Judge Calibration Details
To improve transparency, we describe the calibration protocol of our LLM judge and summarize the associated dataset statistics below.
The calibration set contains 97 search rounds sampled from successful GRPO trajectories with diverse trajectory lengths. Each search round is annotated by a human annotator with respect to two criteria: retrieval utility and reasoning correctness. In addition, the annotator provides a confidence score indicating the certainty of the judgment for each labeled round.
To analyze how judgment quality varies with trajectory length, Table 6 reports the distribution of search rounds across different trajectory lengths, distinguishing whether each round simultaneously satisfies both retrieval utility and reasoning correctness. This allows us to examine how the quality of search rounds evolves as trajectories become longer.
To further understand the uncertainty structure of the calibration data, Table 6 breaks down the annotation confidence conditioned on whether a search round satisfies both quality criteria.
Overall, among the 89 high-confidence rounds, the LLM judge agrees with the human annotator on 87 rounds; among the remaining 8 medium-confidence cases, agreement occurs in 6 cases, yielding an overall agreement rate of 95.8%. This demonstrates strong alignment between the LLM judge and human judgments, particularly in high-confidence annotations.
Number of Search Rounds Satisfies Both Criteria Does not Satisfy
1 5 0
2 17 3
3 18 12
$\geq$4 18 24
Table 5: Quality distribution of search rounds across different trajectory lengths.
Confidence High Medium Low
Satisfies both criteria 55 3 0
Does not satisfy 34 5 0
Table 6: Confidence distribution across quality labels.
Appendix C Computational Cost of LLM Judge
We provide a brief analysis of the computational cost introduced by the LLM judge during training. The judge model GPT-oss-120B has 116.8B parameters and adopts a Mixture-of-Experts (MoE) architecture, where only 5.1B parameters are activated per token. According to official deployment guidelines, the model can be served on a single NVIDIA H100 GPU, making it feasible to integrate into the training pipeline. In practice, incorporating the LLM judge leads to a moderate increase in training cost. Specifically, CW-GRPO training takes 11.7 hours on 5 GPUs (including one GPU dedicated to the judge), compared to 8.8 hours on 4 GPUs for standard GRPO, corresponding to approximately a 33% increase in wall-clock time. Overall, while the LLM judge introduces additional computational cost, we find it manageable in our experimental setup.
Appendix D Prompts
Figure 5 illustrates the prompt used by the LLM judge, which specifies the rubric for assessing each search round in terms of retrieval utility and reasoning correctness, producing the binary contribution signals used in CW-GRPO.
Figure 6 shows the input prompt used during model training, defining how the search agent structures its reasoning, interacts with the retrieval tool across rounds, and generates the final answer.
Figure 5: The prompt for LLM Judge to specify the rubric for assessing each search round in terms of retrieval utility and reasoning correctness.
Figure 6: The prompt for RL training.